RESONANT TWO-PHOTON IONIZATION SPECTROSCOPY OF JET-COOLED OsC

\underline{O.KRECHKIVSKA}, M. D. MORSE, Department of Chemistry, University of Utah, Salt Lake City, Utah, 84112.

The optical spectrum of diatomic OsC has been investigated for the first time, with transitions recorded in the range from 17, 390 to 22, 990 cm^{−1}. The ground state was found to be X^{3}Δ_{3}, deriving from the 4δ^{3}16σ^{1} electronic configuration. Six bands were rotationally resolved and analyzed to obtain ground and excited state rotational constants and bond lengths. Spectra for six OsC isotopomers, ^{192}Os^{12}C (40.3%), ^{190}Os^{12}C (26%), ^{189}Os^{12}C (16%), ^{188}Os^{12}C (13.1%), 187Os^{12}C (1.9%) and ^{186}Os^{12}C (1.6%), were recorded and rotationally analyzed. Four bands were found to originate from the X^{3}Δ_{3} ground state, giving B0^{′′}= 0.533492(33) cm^{−1} and r0^{′′}= 1.67267(5) Å for the ^{192}Os^{12}C isotopomer (1σ error limits); two of these the 0-0 [19.1] 2 ← X^{3}Δ_{3} and 1-0 [19.1] 2 ←X^{3}Δ_{3} bands, form a vibrational progression with ΔG1/2^{′}= 953.019 cm^{−1}. The remaining two bands were identified as originating from an Ω′′= 0 level that remains populated in the supersonic expansion. We believe that this level corresponds to the low-lying A ^{3}Σ_{0+}^{−} state, which derives from the 4δ^{2}16σ^{2} electronic configuration. The OsC molecule differs from the isovalent RuC molecule in having an X^{3}Δ_{3} ground state, rather than the X^{1}Σ^{+} ground state found in RuC. This difference in electronic structure is due to the relativistic stabilization of the 6s orbital in Os, an effect which favors occupation of the 6s−like 16σ orbital. The relativistic stabilization also lowers the energy of the 4δ^{2}16σ^{2}, ^{3}Σ^{−} term, allowing this term to remain populated in the supersonically cooled molecular beam.